# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ III-D WF Defenses
The development of effective defense mechanisms against website fingerprinting has become a critical area of research in online privacy. As introduced in the previous section, WF is a traffic analysis attack that enables an adversary to infer which website a user is visiting by analyzing patterns in encrypted traffic. Despite the use of strong encryption protocols, features such as packet sizes, timing, direction, and burst patterns can still reveal identifiable characteristics of web traffic. When Hintz [83] raised the first WF attack, they also put forward the methods to defend against the WF attack, including adding extra noise data to the traffic, reducing the number of files transferred, and transferring everything in one connection.
@@
Website fingerprinting defenses aim to obscure traffic patterns that attackers rely on to infer which websites users are visiting, even when the data itself is encrypted. Since early research focused on traditional single-proxy systems, similar to defense research, such as link padding [104], traffic morphing [105, 17], and HTTPOS [106]. In 2011, Panchenko et al. [4] made the first defense against WF on Tor. They applied a camouflage strategy to confuse the attack classifiers, which attempted to load another page in parallel to obfuscate the page. Early defenses in Tor primarily focused on padding-based approaches, which insert dummy packets to obscure packet size and burst features. While effective in reducing information leakage, padding often incurs high bandwidth overhead. Later, traffic shaping and regulation techniques were introduced to create more uniform traffic flows by delaying, batching, or smoothing packet transmissions. These methods improve robustness but may add significant latency.
@@
More recent defenses explore traffic morphing strategies, which transform traffic distributions to resemble those of other websites [24, 19]. Additionally, they employ adversarial perturbation techniques inspired by adversarial machine learning [46, 47, 48], which subtly modify traffic patterns to mislead classifiers. Despite these advances, designing an effective WF defense remains challenging due to the trade-off between privacy, usability, and system performance. Many defenses successfully disrupt fine-grained features (e.g., packet order or timing sequences) but remain vulnerable to coarse-grained features such as total timespan, bandwidth ratios, or burst-level statistics. This motivates ongoing research into more comprehensive and adaptive defense mechanisms that can protect against increasingly sophisticated WF attacks. A detailed comparison of the strengths and limitations of representative WF defense strategies will be presented in Section V-B. A brief overview of the WF defense categories in this paper is presented below:
@@
  • Adaptive Padding [53] is a WF defense technique that seeks to hide traffic patterns by inserting dummy packets in a context-aware manner. Unlike fixed padding schemes, adaptive padding selectively adds cover traffic based on observed packet sequences, aiming to obscure identifiable bursts without introducing excessive overhead. By adaptively adjusting padding according to traffic dynamics, adaptive padding reduces the distinctiveness of packet direction and timing features while maintaining lower bandwidth and latency costs compared to constant-rate padding methods.
@@
  • Traffic regularization [107] reshapes network traffic to create a more uniform and predictable flow. Instead of allowing natural bursty patterns, packets are delayed, batched, or transmitted at fixed intervals, making it harder for attackers to exploit timing or burst features. By smoothing out traffic variations, traffic regularization conceals fine-grained patterns, though it often introduces additional latency and may reduce network efficiency.
@@
  • Traffic morphing [9] is a WF defense technique that transforms a website’s traffic patterns to mimic those of another site. By modifying features such as packet sizes, burst lengths, or overall traffic distribution, morphing reduces the distinguishability of traffic traces. This approach aims to confuse classifiers by making different websites appear statistically similar, though it can introduce bandwidth overhead and may not scale well across diverse traffic types.
@@
  • Adversarial perturbation [108] introduces carefully crafted, small modifications to traffic patterns, such as slight delays, dummy packets, or reordering, that are imperceptible to users but mislead ML classifiers. The goal is to degrade the accuracy of WF attacks while keeping bandwidth and latency overhead relatively low. However, its effectiveness depends on the attack model, and adaptive adversaries may still retrain to counter it.
@@ IV-A Dataset Categories
A critical aspect of evaluating WF attacks and defenses is the dataset used for training and testing. These datasets are typically categorized into two main types: closed-world and open-world [9, 11, 94]. The choice of dataset has a significant impact on the attack model’s assumptions and the practical relevance of evaluations.
@@ IV-A1 Closed-world dataset
A closed-world dataset simulates a scenario where an adversary has a complete list of all possible websites a user might visit. The client is assumed to visit a limited set of websites referred to as monitored websites [7, 37, 36]. The dataset consists of a finite, pre-defined set of websites, and the machine learning model is trained to classify a given traffic trace into one of these specific sites. However, this model assumes the adversary has complete knowledge of the user’s browsing space. While this assumption is not realistic for the general internet, it serves as a valuable benchmark for evaluating the maximum potential accuracy of a fingerprinting attack under ideal conditions. Researchers use closed-world datasets to test the effectiveness of new algorithms and features without the added complexity of unknown traffic [95].
@@ IV-A2 Open-world dataset
An open-world dataset provides a more realistic and challenging scenario. In an open-world setting, the dataset is divided into two categories: a small set of ”monitored” or “sensitive” websites and a much larger set of “unmonitored” or “unknown” websites [9]. The attacker’s goal is to accurately identify traffic from the monitored sites while correctly classifying all other traffic as “unknown” [35].
@@
DS-19 [57] 2020 $\times100$ $00×\times1$
@@
ARES [109] 2023 $\times1,00$0 $\times1,00$0
@@ IV-B Public Dataset
Cai et al. [7] collected traffic traces from the Alexa [110] Top 1,000 websites in 2012, ultimately retaining the top 800 cleaned URLs after removing failed pages and resolving redirects. For each URL, they gathered either 20 or 40 traces. Two closed-world Tor traffic datasets were then constructed: (1) one with 100 websites and 20 traces per site, where all packet size information was removed; and (2) another with 100 websites and 40 traces per site, where both packet size and packet direction information were removed. However, no open-world dataset was included. Later, Wang and Goldberg [35] extended this line of work by adding 860 non-monitored traffic instances and 40 monitored instances from Alexa’s Top 1,000 sites to evaluate their model. To collect Tor traffic data, they disabled browser caching, prevented caching to disk, and cleared the browser cache each time the Tor Browser was restarted. However, multi-tab browsing was still not considered, and the collected dataset has limited diversity. Subsequently, Wang et al. [9] released a new dataset in 2014 (named Wang14) that included both closed- and open-world scenarios. They collected 90 instances for each of 100 monitored (sensitive) pages, along with one instance each for 5,000 non-monitored pages from Alexa’s Top 10,000, excluding any sites already included in the monitored set by domain name. In their data collection, webpage loading was performed with regular circuit resets, disabled caching, and time gaps between multiple loads of the same page.
@@
Previous Tor traffic datasets for WF attacks and defenses often overlooked the diversity of sub-pages (e.g., articles, profiles) by focusing primarily on single pages and operating at too small a scale to reflect the real Internet. To address this limitation, Panchenko et al. [11] constructed larger and more diverse datasets. They discarded duplicate entries but deliberately included different web pages from the same website when available. Query parameters such as session identifiers were removed, as they would prevent reproducible retrievals. Furthermore, all URLs pointing to pure advertisement services were excluded, since these are clearly not the user-intended target pages. By combining multiple source datasets [35, 9], the authors produced two large-scale datasets: one containing 34,580 unique websites (118,884 unique web pages) and another containing 65,409 unique websites (211,148 unique web pages). However, Hayes et al. [36] later showed that the feature set of these datasets, which relies heavily on packet order and counts, is highly vulnerable to even simple WF defenses. To conduct their study, they collected two subsets of monitored web pages from Alexa’s top 20,000 websites: (1) 100 instances each from the top 55 Alexa-ranked web pages and (2) 80 instances each from 30 popular Tor hidden services (HS). A Tor hidden service is a website hosted behind a Tor client’s Onion Proxy, which acts as the interface between the application and the network.
@@
Previously collected datasets were relatively limited in size, both in terms of classes (i.e., the number of unique websites) and instances (i.e., the number of traffic traces per website). To address this limitation, Rimmer et al. [94] constructed a closed-world dataset consisting of 900 websites, with 2,500 traffic traces collected for each. They also built an open-world dataset comprising 400,000 unmonitored websites and 200 monitored websites selected from Alexa’s top 400,000 list. To account for temporal variation, Oh et al. [111] collected 110 traffic instances for each of 100 websites across four separate batches from Alexa’s top 100 websites. The batches were collected sequentially, with each subsequent batch starting one week after the previous one. Wang et al. [45] collected their dataset from long-standing pages, selecting 100 of the top sites as the non-sensitive set (after removing duplicates caused by different localizations) and using the next 10,000 pages from Alexa’s top list as the sensitive set. Sirinam et al. [93] built a closed-world dataset of 95 websites with 1,000 traces per site from Alexa’s top 100 list, as well as an open-world dataset containing 40,716 traffic traces.
@@
In 2020, Gong and Wang [57] collected a dataset, referred to as DS-19, from the homepages of Alexa’s top 100 websites. They gathered 100 traces of each homepage as the monitored set and 10,000 additional web pages as the unmonitored set, filtering out pages that failed to load (e.g., those inaccessible via Tor). Recognizing that an attacker cannot rely on training datasets containing only correct pairings, Rahman et al. [12] collected Tor traffic from 100 monitored websites and sampled 10,000 unmonitored websites from Alexa’s top 14,000 list. To accurately represent the browsing behavior of the Tor Browser bundle, the authors used a modified variant of the Tor Browser Crawler [112]. Similarly, Cadena et al. [113] built an open-world dataset consisting of 11,307 popular Alexa websites (excluding the top 100) along with 100 traces for each of Alexa’s top 100 websites. Wang [114] selected the top 100 Alexa pages as the sensitive set, visiting each 200 times, and the next 80,000 pages as the non-sensitive set, visiting each once. Gong et al. [115] created a new undefended dataset using WFDefProxy [116], consisting of Tor traces from 100 monitored pages and 10,000 non-monitored pages, where each monitored page was visited 100 times and each non-monitored page once. To study the impact of concept drift in WF attacks, Bahramali et al. [71] collected up to 550 traces per visit for each of 225 non-onion websites in the closed-world setting. For the open-world setting, the authors gathered a single trace per website, selecting 10,000 unmonitored websites. The collected data was then divided into three subsets: (1) Drift90: a set of 90 monitored websites, each with at least 100 superior and 20 inferior traces; (2) Drift-Guard: the same 90 monitored websites, further split into traces collected through 11 guard relays in Europe and 7 guard relays in North America; and (3) Drift5000: a set of 5,000 unmonitored websites.
@@
To evaluate their proposed WF attack model in a multi-tab setting, Xu et al. [62] collected both single-tab and two-tab datasets. The single-tab dataset consists of three subsets: (i) 50 instances from each of 50 monitored web pages without background noise for training, (ii) another 50 instances from each of the same 50 monitored web pages without background noise for testing, and (iii) an open-world subset comprising instances from 2,500 unmonitored web pages. The two-tab dataset was constructed with five different time gaps: two, three, four, five, and six seconds. For each time gap, they collected traffic traces from 50 monitored web pages, with 50 instances per page. In addition, they collected 5,000 instances for use in dynamic split experiments. Guan et al. [117] collected 10,000 training traces and 1,000 testing traces from 50 websites. They generated real-world two-tab traces by visiting two sites 100 times each and combining the $\times100$ possible pairs. They also constructed synthetic two- and three-tab traces by merging single-tab traces from the dataset in [45], selecting the 50 websites with the most packet traces (90 traces per site for training, the rest for testing). Each two-tab trace was formed by aligning two single-tab traces with an initial delay and interleaving packets by timestamp, ensuring every site appeared in both tab positions at least once, while the other tab was chosen randomly. To better reflect real-world scenarios where multiple tabs may overlap, they further merged three traces per sample in the closed-world setting, with the additional two tabs randomly chosen from the 50 websites. This resulted in 13,500 training traces and 2,562 testing traces. Deng et al. [109] extended this line of work by collecting a five-tab dataset that comprises 10 dark web websites and 90 websites outside Alexa’s top 100 list. They also investigated scenarios with multiple Tor Browser versions, gathering five-tab traces from versions 10.0.15, 10.5, 11.0.3, 11.0.10, and 11.5. In addition, they built a five-tab dataset from 1,000 subpages of 100 websites.
@@
Existing multi-tab datasets are typically created by manually synthesizing multi-tab traces from single-tab recordings. To enable a more realistic evaluation of multi-tab WF attacks, Jin et al. [43] collected both single-tab and 2-tab datasets for Chrome and Tor browsers (TBB). Their monitored set comprises 5,000 traces (50 monitored websites × 100 single-tab traces per site), and they collected an additional 5,000 non-monitored traces for each browser. For the open-world evaluation, they gathered over 10,000 2-tab trace samples for Chrome and Tor; these traces were generated by randomly selecting overlap ratios in the range [0.1, 0.5] so as to simulate diverse access and overlap scenarios. For traces with more than two tabs, the authors continued to rely on manual synthesis: multi-tab samples are created by concatenating single-tab traces using a controlled overlap ratio or by introducing randomized inter-tab delays to emulate real browsing timing. These synthesis strategies allow the authors to scale training sets, vary overlap characteristics, and evaluate model robustness across a range of realistic and adversarial multi-tab conditions.

### v2

@@ III-D WF Defenses
The development of effective defense mechanisms against website fingerprinting has become a critical area of research in online privacy. As introduced in the previous section, WF is a traffic analysis attack that enables an adversary to infer which website a user is visiting by analyzing patterns in encrypted traffic. Despite the use of strong encryption protocols, features such as packet sizes, timing, direction, and burst patterns can still reveal identifiable characteristics of web traffic. When Hintz [hintz2002fingerprinting] raised the first WF attack, they also put forward the methods to defend against the WF attack, including adding extra noise data to the traffic, reducing the number of files transferred, and transferring everything in one connection.
@@
Website fingerprinting defenses aim to obscure traffic patterns that attackers rely on to infer which websites users are visiting, even when the data itself is encrypted. Since early research focused on traditional single-proxy systems, similar to defense research, such as link padding [fu2003analytical], traffic morphing [wright2009traffic, nithyanand2014glove], and HTTPOS [luo2011httpos]. In 2011, Panchenko et al. [panchenko2011website] made the first defense against WF on Tor. They applied a camouflage strategy to confuse the attack classifiers, which attempted to load another page in parallel to obfuscate the page. Early defenses in Tor primarily focused on padding-based approaches, which insert dummy packets to obscure packet size and burst features. While effective in reducing information leakage, padding often incurs high bandwidth overhead. Later, traffic shaping and regulation techniques were introduced to create more uniform traffic flows by delaying, batching, or smoothing packet transmissions. These methods improve robustness but may add significant latency.
@@
More recent defenses explore traffic morphing strategies, which transform traffic distributions to resemble those of other websites [ling2024wfguard, shen2024real]. Additionally, they employ adversarial perturbation techniques inspired by adversarial machine learning [gong2022surakav, qiao2024trace, jiang2024rudolf], which subtly modify traffic patterns to mislead classifiers. Despite these advances, designing an effective WF defense remains challenging due to the trade-off between privacy, usability, and system performance. Many defenses successfully disrupt fine-grained features (e.g., packet order or timing sequences) but remain vulnerable to coarse-grained features such as total timespan, bandwidth ratios, or burst-level statistics. This motivates ongoing research into more comprehensive and adaptive defense mechanisms that can protect against increasingly sophisticated WF attacks. A detailed comparison of the strengths and limitations of representative WF defense strategies will be presented in Section V-B. A brief overview of the WF defense categories in this paper is presented below:
@@
  • Adaptive Padding [shmatikov2006timing] is a WF defense technique that seeks to hide traffic patterns by inserting dummy packets in a context-aware manner. Unlike fixed padding schemes, adaptive padding selectively adds cover traffic based on observed packet sequences, aiming to obscure identifiable bursts without introducing excessive overhead. By adaptively adjusting padding according to traffic dynamics, adaptive padding reduces the distinctiveness of packet direction and timing features while maintaining lower bandwidth and latency costs compared to constant-rate padding methods.
@@
  • Traffic regularization [wang2021one] reshapes network traffic to create a more uniform and predictable flow. Instead of allowing natural bursty patterns, packets are delayed, batched, or transmitted at fixed intervals, making it harder for attackers to exploit timing or burst features. By smoothing out traffic variations, traffic regularization conceals fine-grained patterns, though it often introduces additional latency and may reduce network efficiency.
@@
  • Traffic morphing [wang2014effective] is a WF defense technique that transforms a website’s traffic patterns to mimic those of another site. By modifying features such as packet sizes, burst lengths, or overall traffic distribution, morphing reduces the distinguishability of traffic traces. This approach aims to confuse classifiers by making different websites appear statistically similar, though it can introduce bandwidth overhead and may not scale well across diverse traffic types.
@@
  • Adversarial perturbation [li2019dynamic] introduces carefully crafted, small modifications to traffic patterns, such as slight delays, dummy packets, or reordering, that are imperceptible to users but mislead ML classifiers. The goal is to degrade the accuracy of WF attacks while keeping bandwidth and latency overhead relatively low. However, its effectiveness depends on the attack model, and adaptive adversaries may still retrain to counter it.
@@ IV-A Dataset Categories
A critical aspect of evaluating WF attacks and defenses is the dataset used for training and testing. These datasets are typically categorized into two main types: closed-world and open-world [wang2014effective, panchenko2016website, rimmer2017automated]. The choice of dataset has a significant impact on the attack model’s assumptions and the practical relevance of evaluations.
@@ IV-A1 Closed-world dataset
A closed-world dataset simulates a scenario where an adversary has a complete list of all possible websites a user might visit. The client is assumed to visit a limited set of websites referred to as monitored websites [cai2012touching, wang2016realistically, hayes2016k]. The dataset consists of a finite, pre-defined set of websites, and the machine learning model is trained to classify a given traffic trace into one of these specific sites. However, this model assumes the adversary has complete knowledge of the user’s browsing space. While this assumption is not realistic for the general internet, it serves as a valuable benchmark for evaluating the maximum potential accuracy of a fingerprinting attack under ideal conditions. Researchers use closed-world datasets to test the effectiveness of new algorithms and features without the added complexity of unknown traffic [sirinam2019triplet].
@@ IV-A2 Open-world dataset
An open-world dataset provides a more realistic and challenging scenario. In an open-world setting, the dataset is divided into two categories: a small set of ”monitored” or “sensitive” websites and a much larger set of “unmonitored” or “unknown” websites [wang2014effective]. The attacker’s goal is to accurately identify traffic from the monitored sites while correctly classifying all other traffic as “unknown” [wang2013improved].
@@
DS-19 [gong2020zero] 2020 $\times100$ $00×\times1$
@@
ARES [deng2023robust] 2023 $\times1,00$0 $\times1,00$0
@@ IV-B Public Dataset
Cai et al. [cai2012touching] collected traffic traces from the Alexa [alexa2025top] Top 1,000 websites in 2012, ultimately retaining the top 800 cleaned URLs after removing failed pages and resolving redirects. For each URL, they gathered either 20 or 40 traces. Two closed-world Tor traffic datasets were then constructed: (1) one with 100 websites and 20 traces per site, where all packet size information was removed; and (2) another with 100 websites and 40 traces per site, where both packet size and packet direction information were removed. However, no open-world dataset was included. Later, Wang and Goldberg [wang2013improved] extended this line of work by adding 860 non-monitored traffic instances and 40 monitored instances from Alexa’s Top 1,000 sites to evaluate their model. To collect Tor traffic data, they disabled browser caching, prevented caching to disk, and cleared the browser cache each time the Tor Browser was restarted. However, multi-tab browsing was still not considered, and the collected dataset has limited diversity. Subsequently, Wang et al. [wang2014effective] released a new dataset in 2014 (named Wang14) that included both closed- and open-world scenarios. They collected 90 instances for each of 100 monitored (sensitive) pages, along with one instance each for 5,000 non-monitored pages from Alexa’s Top 10,000, excluding any sites already included in the monitored set by domain name. In their data collection, webpage loading was performed with regular circuit resets, disabled caching, and time gaps between multiple loads of the same page.
@@
Previous Tor traffic datasets for WF attacks and defenses often overlooked the diversity of sub-pages (e.g., articles, profiles) by focusing primarily on single pages and operating at too small a scale to reflect the real Internet. To address this limitation, Panchenko et al. [panchenko2016website] constructed larger and more diverse datasets. They discarded duplicate entries but deliberately included different web pages from the same website when available. Query parameters such as session identifiers were removed, as they would prevent reproducible retrievals. Furthermore, all URLs pointing to pure advertisement services were excluded, since these are clearly not the user-intended target pages. By combining multiple source datasets [wang2013improved, wang2014effective], the authors produced two large-scale datasets: one containing 34,580 unique websites (118,884 unique web pages) and another containing 65,409 unique websites (211,148 unique web pages). However, Hayes et al. [hayes2016k] later showed that the feature set of these datasets, which relies heavily on packet order and counts, is highly vulnerable to even simple WF defenses. To conduct their study, they collected two subsets of monitored web pages from Alexa’s top 20,000 websites: (1) 100 instances each from the top 55 Alexa-ranked web pages and (2) 80 instances each from 30 popular Tor hidden services (HS). A Tor hidden service is a website hosted behind a Tor client’s Onion Proxy, which acts as the interface between the application and the network.
@@
Previously collected datasets were relatively limited in size, both in terms of classes (i.e., the number of unique websites) and instances (i.e., the number of traffic traces per website). To address this limitation, Rimmer et al. [rimmer2017automated] constructed a closed-world dataset consisting of 900 websites, with 2,500 traffic traces collected for each. They also built an open-world dataset comprising 400,000 unmonitored websites and 200 monitored websites selected from Alexa’s top 400,000 list. To account for temporal variation, Oh et al. [oh2017p] collected 110 traffic instances for each of 100 websites across four separate batches from Alexa’s top 100 websites. The batches were collected sequentially, with each subsequent batch starting one week after the previous one. Wang et al. [wang2017walkie] collected their dataset from long-standing pages, selecting 100 of the top sites as the non-sensitive set (after removing duplicates caused by different localizations) and using the next 10,000 pages from Alexa’s top list as the sensitive set. Sirinam et al. [sirinam2018deep] built a closed-world dataset of 95 websites with 1,000 traces per site from Alexa’s top 100 list, as well as an open-world dataset containing 40,716 traffic traces.
@@
In 2020, Gong and Wang [gong2020zero] collected a dataset, referred to as DS-19, from the homepages of Alexa’s top 100 websites. They gathered 100 traces of each homepage as the monitored set and 10,000 additional web pages as the unmonitored set, filtering out pages that failed to load (e.g., those inaccessible via Tor). Recognizing that an attacker cannot rely on training datasets containing only correct pairings, Rahman et al. [rahman2019tik] collected Tor traffic from 100 monitored websites and sampled 10,000 unmonitored websites from Alexa’s top 14,000 list. To accurately represent the browsing behavior of the Tor Browser bundle, the authors used a modified variant of the Tor Browser Crawler [mjuarezm2021webfp]. Similarly, Cadena et al. [de2020trafficsliver] built an open-world dataset consisting of 11,307 popular Alexa websites (excluding the top 100) along with 100 traces for each of Alexa’s top 100 websites. Wang [wang2020high] selected the top 100 Alexa pages as the sensitive set, visiting each 200 times, and the next 80,000 pages as the non-sensitive set, visiting each once. Gong et al. [gong2024wfcat] created a new undefended dataset using WFDefProxy [gong2023wfdefproxy], consisting of Tor traces from 100 monitored pages and 10,000 non-monitored pages, where each monitored page was visited 100 times and each non-monitored page once. To study the impact of concept drift in WF attacks, Bahramali et al. [bahramali2023realistic] collected up to 550 traces per visit for each of 225 non-onion websites in the closed-world setting. For the open-world setting, the authors gathered a single trace per website, selecting 10,000 unmonitored websites. The collected data was then divided into three subsets: (1) Drift90: a set of 90 monitored websites, each with at least 100 superior and 20 inferior traces; (2) Drift-Guard: the same 90 monitored websites, further split into traces collected through 11 guard relays in Europe and 7 guard relays in North America; and (3) Drift5000: a set of 5,000 unmonitored websites.
@@
To evaluate their proposed WF attack model in a multi-tab setting, Xu et al. [xu2018multi] collected both single-tab and two-tab datasets. The single-tab dataset consists of three subsets: (i) 50 instances from each of 50 monitored web pages without background noise for training, (ii) another 50 instances from each of the same 50 monitored web pages without background noise for testing, and (iii) an open-world subset comprising instances from 2,500 unmonitored web pages. The two-tab dataset was constructed with five different time gaps: two, three, four, five, and six seconds. For each time gap, they collected traffic traces from 50 monitored web pages, with 50 instances per page. In addition, they collected 5,000 instances for use in dynamic split experiments. Guan et al. [guan2021bapm] collected 10,000 training traces and 1,000 testing traces from 50 websites. They generated real-world two-tab traces by visiting two sites 100 times each and combining the $\times100$ possible pairs. They also constructed synthetic two- and three-tab traces by merging single-tab traces from the dataset in [wang2017walkie], selecting the 50 websites with the most packet traces (90 traces per site for training, the rest for testing). Each two-tab trace was formed by aligning two single-tab traces with an initial delay and interleaving packets by timestamp, ensuring every site appeared in both tab positions at least once, while the other tab was chosen randomly. To better reflect real-world scenarios where multiple tabs may overlap, they further merged three traces per sample in the closed-world setting, with the additional two tabs randomly chosen from the 50 websites. This resulted in 13,500 training traces and 2,562 testing traces. Deng et al. [deng2023robust] extended this line of work by collecting a five-tab dataset that comprises 10 dark web websites and 90 websites outside Alexa’s top 100 list. They also investigated scenarios with multiple Tor Browser versions, gathering five-tab traces from versions 10.0.15, 10.5, 11.0.3, 11.0.10, and 11.5. In addition, they built a five-tab dataset from 1,000 subpages of 100 websites.
@@
Existing multi-tab datasets are typically created by manually synthesizing multi-tab traces from single-tab recordings. To enable a more realistic evaluation of multi-tab WF attacks, Jin et al. [jin2023transformer] collected both single-tab and 2-tab datasets for Chrome and Tor browsers (TBB). Their monitored set comprises 5,000 traces (50 monitored websites × 100 single-tab traces per site), and they collected an additional 5,000 non-monitored traces for each browser. For the open-world evaluation, they gathered over 10,000 2-tab trace samples for Chrome and Tor; these traces were generated by randomly selecting overlap ratios in the range [0.1, 0.5] so as to simulate diverse access and overlap scenarios. For traces with more than two tabs, the authors continued to rely on manual synthesis: multi-tab samples are created by concatenating single-tab traces using a controlled overlap ratio or by introducing randomized inter-tab delays to emulate real browsing timing. These synthesis strategies allow the authors to scale training sets, vary overlap characteristics, and evaluate model robustness across a range of realistic and adversarial multi-tab conditions.